OpenStudy (anonymous):

are the pilots actions good or bad in the cold equation what examples from the story make you think this

OpenStudy (anonymous):

man i need help son

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) (01.07) Solve 3x - x + 2 = 12. -5 4 7 5 -------------------------------------------------------------------------------- Question 2 (Multiple Choice Worth 1 points) (01.07) Name the algebraic property demonstrated in the example below: 2 + x + y = x + 2 + y Commutative Property of Addition Associative Property of Addition Reflexive Property Transitive Property -------------------------------------------------------------------------------- Question 3 (Multiple Choice Worth 1 points) (01.07) Solve for x: = 2. x = -2 x = 2 x = x = -------------------------------------------------------------------------------- Question 4 (Multiple Choice Worth 1 points) (01.07) Solve + 5 = 13. -16 16 -36 36 -------------------------------------------------------------------------------- Question 5 (Multiple Choice Worth 1 points) (01.07) Emma took 35 pictures of flowers on a nature walk and an additional 3 pictures of each bird she saw. She took a total of 83 pictures. Which equation can be used to find the number of birds, x, that Emma saw on her walk? 35x + 3x = 83 35 + 3 = 83x 35x + 3 = 83 35 + 3x = 83 -------------------------------------------------------------------------------- Question 6 (Multiple Choice Worth 1 points) (01.07) Solve 3c + 4 = -20. -5.3 6 8 -8 -------------------------------------------------------------------------------- Question 7 (Multiple Choice Worth 1 points) (01.07) Solve for x: (2x − 14) = 4 15 9 11 -------------------------------------------------------------------------------- Question 8 (Multiple Choice Worth 1 points) (01.07) Solve 4(x + 1) + 8 = 24. 3.75 -2 5 3 -------------------------------------------------------------------------------- Question 9 (Multiple Choice Worth 1 points) (01.07) Name the algebraic property demonstrated in the example below: 3 • (x • y) = (3 • x) • y Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication -------------------------------------------------------------------------------- Question 10 (Multiple Choice Worth 1 points) (01.07) Which of the following describes the correct process for solving the equation 3x + 6 = 21 and arrives at the correct solution? Divide both sides by 6 and then add 3. The solution is x = 5. Add 6 to both sides of the equation and then divide by 3. The solution is x = 9. Subtract 6 from both sides of the equation and then divide by 3. The solution is x = 8. Subtract 6 from both sides of the equation and then divide by 3. The solution is x = 5.